OpenStudy (anonymous):

The volume of a sphere is 5,000pi m^3. What is the surface area of the sphere to the nearest square meter?

OpenStudy (campbell_st):

Volume of the Sphere you need to find the radius \[5000pi = \frac{4}{3} \pi r^3\] \[r^3 = 3750 \] \[r = \sqrt[3]{3750} = 15.5632\] Surface Area \[SA = 4 \pi (15.5632)^2\] I'll leave you to calculate the surface area

OpenStudy (anonymous):

Thank you so much!! :) I really appriciate it!!

OpenStudy (campbell_st):

hope it helps